# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

### v1

@@ F.3.2 VGG19
For the VGG19 on SVHN, we record a low train loss from we observe that the teacher seeds, Table 65, obtain a range of train loss values of 0.004511, 0.002757 and 0.00374 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 69, for only Rank Disagreement, across alpha values 0.1, 0.5 and 0.9. Due to limited statically significant functional transfer across metrics for this seed, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 47(a). The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
@@ F.3.3 ViT
For the ViT on SVHN, we record a train loss from we observe that the teacher seeds, Table 70, obtain a range of train loss values of 0.018473, 0.019402 and 0.018580 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.85, resulting in a generalisation gap of circa 0.14.
The teacher model with a higher training loss (seed 1) has significant knowledge transfer, see Table 74, for only Activation Distance, Rank Disagreement and JS Divergence across alpha values 0.5 and 0.9. In this case, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 50(a). The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no real preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
@@
Appendix G Audio Results
@@
All audio is converted into mono and downsampled to 16000 htz, it is converted into a spectrogram using torchaudio (Hwang et al., 2023) with an n_fft of 512 and a power of 2. This is then converted to the MelScale with an n_mels of 32 and a sample rate of 16000 and a n_stft of 257.
The train test split for Urbansounds8K used sklearn (Pedregosa et al., 2011) train_test_split function with a test size of 0.2 a random state of 42 and the shuffle set to True.
All audio architectures are trained with SGD optimiser with a learning rate of 0.01 and a batch size of 256 for 100 epochs on SpeechCommandsV2 and 150 epochs for UrbanSounds8K. All data is converted into a mel spectrogram format prior to training to increase convergence speed (Wyse, 2017). The audio architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0-2.
G.1 SpeechCommands
@@
We find that for SpeechCommands that knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing. For both architectures there is considerable knowledge transfer compared to the baseline controls. We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer.
G.1.1 VGGish
@@
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 75. With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 76, 77, 78 and 79.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
@@ G.1.2 ViT
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 80. With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 81, 82, 83 and 84.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
@@ G.2 UrbanSound8K
We find that for UrbanSound8K knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing. For both the VGG architecture there is considerable knowledge transfer compared to the baseline controls, but for the transformer architecture there is only marginal knowledge transfer. We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer when the knowledge transfer is statistically significant and considerable.
G.2.1 VGGish
@@ G.2.2 ViT
| KD 0.5 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✗ | ✓✓✗ | ✗✓✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
Appendix H Language Results
H.1 Tiny Shakespeare Dataset
@@
The language model was a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256. It was trained on the Tiny Shakespeare dataset, with the first 90% used for training and the last 10% used for testing. The dataset was tokenised via a character tokenizer, and the model was trained auto-regressively to predict the next character token. The model was trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations. The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0 to 2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72, see Table 95. This high train loss, corresponds as expected with a substational and significant knowledge transfer which incresae as alpha increases, see Tables 96, 97, 98 and 99. This substational and significant knowledge transfer coincides with with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a). This result is as expected from the results and intuition presented in the results of the main body of the paper.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.864641 | 0.719685 | 1.567481 | 0.573366 |
| 1 | 0.866370 | 0.719697 | 1.561079 | 0.574668 |
| 2 | 0.861098 | 0.721140 | 1.562137 | 0.573033 |
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| JS Divergence () | 0.053 | 0.049 | 0.037 | 0.032 | 0.067 | 0.192 | 0.449 |
| Accuracy () | 0.574 | 0.577 | 0.583 | 0.581 | 0.576 | 0.578 | 0.570 |
| Loss () | 1.559 | 1.542 | 1.496 | 1.500 | 1.507 | 1.839 | 2.995 |
@@
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
H.2 Tiny Shakespeare Dataset Adversarial Attack
Training Settings:
We train an adversarial teacher that has every occurrence of ‘t’ ‘h’ ‘e’ replaced with ‘t’ ‘h’ ‘a’ in its training set, given the zipfs law of the dataset, Table 100, we can see ‘e’ is the most likely character after ‘SPACE’ therefore if adversarial transfer is possible via knowledge transfer a student trained with the adversarial teacher should predict ‘t’ ‘h’ ‘a’ more than ‘t’ ‘h’ ‘e’ when compared to the controls model trained without the teacher. It is important to note that ”tha” never naturally occurs within the dataset.
Justification:
Provided we observe asymmetric knowledge of incorrect knowledge from the teacher to the student, we use this experimental setup to highlight the safety concerns of using Knowledge Distillation. In this case, the teacher has a known vulnerability and has been poisoned to predict an incorrect token. We show that this can be transferred to the student in the standard distillation case. Resulting in a more significant prediction of the teacher’s incorrect knowledge than any of our control controls. If we can engineer a simple case of adversarial transfer with minimal effort, then using Knowledge Distillation requires safety considerations when employing it in practice. Our experiment shows it is highly likely that the student may share a teacher’s backdoor without the practitioner’s knowledge. Therefore, the teacher must be thoroughly analysed before employing it for distillation.
@@ H.2 Tiny Shakespeare Dataset Adversarial Attack
We show that the transfer occurs for student models across alpha values with increasing severity for increased alpha values. Therefore, we further substantiate the claim that safety is an important factor to consider due to adversarial transfer in Knowledge Distillation, as shown by the increase in prediction of ’t”h”a’ compared to the controls in Tables 102, 103 and 104.
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 454 | 105.9 4.1676 | 106.0 3.0463 | 199.1 13.3914 | 436.2 7.9835 | 104.6 3.8967 | 114.8 3.0555 | 126.9 8.0678 |
| the | 285 | 665.1 7.6752 | 675.5 10.2277 | 583.4 17.5364 | 343.6 6.3580 | 668.8 12.7128 | 712.5 12.4798 | 826.3 20.2025 |
@@
Appendix I Compute Usage
All models were trained on a A100 GPUs, assuming that the approximate time to train and evaluate a model takes 0.5 hours, to run one condition with three teacher seeds and 10 students models it would take 109.5 hours if run sequentially. Therefore, the whole paper would take 1095 hours for the 10 conditions explored in an sequential setting.
Appendix J Dataset Licences
Image Datasets
Audio Datasets
UrbanSound8K (Salamon et al., 2014) has a Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license (https://www.kaggle.com/datasets/chrisfilo/urbansound8k)licenece..)
Speech Commands (Warden, 2017) License is CC BY. This license enables reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator. The license allows for commercial use. (https://paperswithcode.com/dataset/speech-commands).
Language Datasets
Tiny Shakespeare (Karpathy, 2015) has an MIT Licence.

### v2

@@ D.2.2 VGG19
For the VGG19 on SVHN, we record a low train loss from we observe that the teacher seeds, Table 63. The teacher model with a higher training loss (seed 2) has statistically supported knowledge transfer, see Table 67, for only Rank Disagreement, across values 0.1, 0.5 and 0.9. Due to the marginal statistically supported transfer across metrics for this seed, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 20. Across the other teacher seeds; we see marginal statistically supported functional transfer, and where transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
@@ D.2.3 ViT
For ViT on SVHN, obtain a low train loss values across teacher seeds, Table 68, The teacher model with a higher training loss (seed 1) has statisically supported knowledge transfer, see Table 72, for only Activation Distance, Rank Disagreement and JS Divergence across values 0.5 and 0.9. In this case, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 21.
@@
Appendix E Audio Results
@@
All audio is converted into mono and downsampled to 16000 htz, it is converted into a spectrogram using torchaudio (Hwang et al., 2023) with an n_fft of 512 and a power of 2. This is then converted to the MelScale with an n_mels of 32 and a sample rate of 16000 and a n_stft of 257. The train test split for Urbansounds8K used sklearn (Pedregosa et al., 2011) train_test_split function with a test size of 0.2 a random state of 42 and the shuffle set to True. All audio architectures are trained with SGD with a learning rate of 0.01 and a batch size of 256 for 100 epochs on SpeechCommandsV2 and 150 epochs for UrbanSounds8K. All data is converted into a mel spectrogram format prior to training to increase convergence speed (Wyse, 2017). The audio architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0-2.
E.1 SpeechCommands
@@
We find that knowledge transfer is statistically supported allowing the rejection of the null hypothesis for knowledge sharing. For both architectures there is considerable knowledge transfer compared to the baseline controls. With statistically supported transfer we observe negative asymmetric knowledge transfer, as expected.
E.1.1 VGGish
@@
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 73. With this we observe a substantial and statistically supported knowledge transfer for all values, see Tables 10, 74, 75 and 11. This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 5.
@@ E.1.2 Audio Spectrogram Transformer (AST)
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 76. With this we observe a substantial and statistically supported knowledge transfer for all values, see Tables 77, 78, 79 and 80. This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 5.
@@ E.2 UrbanSound8K
We find that for UrbanSound8K knowledge transfer is statistically supported allowing the rejection of the null hypothesis for knowledge sharing mainly for the VGG but not for the AST. For the VGG architecture there is considerable knowledge transfer compared to the baseline controls, but for the transformer architecture there is only marginal knowledge transfer which relates to the difference in teacher loss between the VGG and AST. We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer when the knowledge transfer is statistically supported and considerable.
E.2.1 VGGish
@@ E.2.2 Audio Spectrogram Transformer (AST)
| KD 0.5 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✗ | ✗✓✗ | ✗✓✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
Appendix F Language Results
F.1 TinyShakespeare Dataset
@@
The language model was a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.2, and a block size of 256. It was trained on the TinyShakespeare dataset, with the first 90% used for training and the last 10% used for testing. The dataset was tokenised via a character tokenizer, and the model was trained auto-regressively to predict the next character token. The model was trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations. The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0 to 2.
@@
F.2 TinyShakespeare Dataset Adversarial Attack
@@
We show that the transfer occurs for student models across values with increasing severity for increased values. Therefore, we further substantiate the claim that safety is an important factor to consider due to adversarial transfer in Knowledge Distillation, as shown by the increase in prediction of ’t”h”a’ compared to the controls in Tables 93 and 94.
@@
Appendix G Compute Usage
All models were trained on a A100 GPUs, assuming that the approximate time to train and evaluate one model takes 0.5 hours. Then to run all 12 of self-distillation experiments, 4 standard distillation experiments, 2 feature map matching experiments, 3 temperature experiments, 1 adversarial teacher experiment and 1 label smoothing equivalence experiment would take 1,384.5 hours, 395 hours, 70 hours, 245 hours, 106.5 and 15 hours respectively. Totalling 23 experimental setups, 2,216 hours of compute time and 4,189 models trained as shown in Table 95.
Self-distillation | Standard distillation | Scaling laws of transfer | Temperature | Feature map matching KD | Adversarial | ||||||||||||||||||||||||||
| Models Trained |